When 0.149 g of lactic acid is combusted in a constant volume calorimeter in an excess of oxygen, only CO2 and water are products and 2.24 kJ of energy is released. If the heat capacity of the calorimeter is 0.827 kJ/K and the initial temperature was 23.44 °C, what is the final temperature in °C?




1
Expert's answer
2021-11-22T17:03:01-0500

Molar mass of lactic acid= 90.08

=0.149/90.08= 0.00165mol

2.24= 0.827×0.00165×(x-296.44)

2.24=0.014× (x-296.44)

2.24=0.014x -0.405

0.014x= 2.24+0.405

0.014x=2.645

X=189°C